OpenStudy (anonymous):

log5(lower) (4x-7)=2 Can anyone solve this one?

OpenStudy (shadowfiend):

So is that: \[log_5(4x - 7) = 2\] ?

OpenStudy (anonymous):

yes!

OpenStudy (shadowfiend):

Ok. Well, first off remember that: \[log_a b = c\] Means \[a^c = b\] So in this case we have: \[5^2 = 4x - 7\] Can you solve that?

OpenStudy (anonymous):

yes, thanx a lot!

OpenStudy (shadowfiend):

No problem :)